# A Framework for Synthetic Audio Conversations Generation using Large Language Models

### v1

@@ I Introduction
The emergence of large language models (LLMs) has revolutionized natural language processing, enabling the generation of coherent and contextually relevant text. These advancements have propelled various applications, from automated content creation to sophisticated dialogue systems, significantly enhancing human-computer interactions. However, the potential of these models extends beyond text generation, particularly in the realm of audio-based applications. One under-explored area is the creation of synthetic audio datasets for multi-speaker conversations, which are critical for training and evaluating models in numerous speech-related applications. However, existing audio conversation datasets are relatively uncommon and tend to be limited in size and diversity, posing challenges in developing robust and adaptable models capable of handling diverse scenarios. To address this gap, we propose ConversaSynth, a comprehensive framework for generating synthetic audio conversations using LLMs and text-to-speech (TTS) systems. Our framework leverages the capabilities of LLMs to generate dialogues involving multiple speakers with distinct personas, covering a wide range of topics. These text-based dialogues are then converted into audio, resulting in a contextually rich and varied synthetic conversation audio dataset. The generated datasets demonstrate substantial diversity and realism, which are crucial for training models to perform accurately in real-world scenarios. By utilizing LLMs to create coherent dialogues and employing TTS models to generate natural sounding speech, ConversaSynth can create a dataset that can serve as a valuable resource for various applications, including:
  • Audio Classification: Improving the accuracy of models tasked with identifying and categorizing different types of audio content.
  • Speech Recognition and Extraction: Advancing the development of models capable of accurately transcribing and segmenting multi-speaker conversations.
Moreover, the flexibility of our framework allows for the creation of specific conversational topics and speaker attributes that are underrepresented in existing datasets. This controlled environment is essential for training models that need to perform reliably across a broad spectrum of contexts and speaker variations. In summary, ConversaSynth represents a significant advancement in the generation of synthetic audio datasets. By bridging the gap between text-based dialogue generation and audio synthesis, our framework not only addresses the current limitations in the field but also creates the way for future research and development in synthetic audio generation and its applications in machine learning and artificial intelligence.
@@ II LITERATURE REVIEW
The field of artificial intelligence (AI) has witnessed transformative advancements in Natural Language Processing (NLP) and Text-to-Speech (TTS) technologies. The integration of these technologies has the potential to revolutionize applications such as virtual assistants, content creation, and accessibility tools. This literature review explores the evolution and advancements in LLMs and TTS systems, and examines existing research on their integration.
@@ II-A Large Language Models (LLMs)
Large Language Models (LLMs) have significantly impacted the field of NLP by utilizing massive datasets to generate human-like text. The progression from models like GPT-2 to GPT-3 and GPT-4 has marked substantial improvements in text generation capabilities [1]. These models excel at understanding context, syntax, and semantics, making them proficient in a range of language tasks. The Transformer architecture, introduced by Vaswani in 2017 [2], has been pivotal in advancing LLMs. This architecture employs self-attention mechanisms that allow models to weigh the importance of each word in a sentence relative to others, thereby capturing long-range dependencies and complex language patterns. With the introduction of the Transformer architecture, the development of models that can generate more coherent and contextually relevant text has emerged. In 2020, OpenAI released GPT-3 with 175 billion parameters, set new benchmarks for text generation. Its ability to generate contextually appropriate and coherent text across various tasks has demonstrated the potential of LLMs in content creation and interactive applications. Later, GPT-4 was released with even more parameters and refined training techniques, showing improved versatility and accuracy in generating and understanding text. With these advancements, LLMs have revolutionized text generation, enabling the creation of high-quality articles, creative content, and automated responses. Their ability to generate human-like text has applications in writing assistance, creative storytelling, and chatbots.
@@ II-B Text-to-Speech (TTS) Technologies
Text-to-Speech (TTS) technology has advanced significantly, evolving from basic methods to sophisticated deep learning techniques. This progress has made TTS systems much better at producing natural and clear speech from text. Early TTS systems used a method called concatenative synthesis. In this method, pre-recorded pieces of speech were joined together to form sentences. While this worked, it often resulted in speech that sounded less natural due to the limited variety and quality of the recorded sounds. Later, the introduction of deep learning models like Tacotron [3] and WaveNet [4] revolutionized TTS. Tacotron converts text into a visual representation of sound called a mel-spectrogram, which is then turned into an audio waveform by a vocoder, such as WaveNet. This end-to-end approach greatly enhanced the quality and realism of synthetic speech. Newer approaches, like Parler-TTS [5] and XTTS [6], have pushed the boundaries even further. Parler-TTS is designed to generate expressive speech, making the output sound more engaging and human-like. XTTS allows the generation of highly customized and context-aware speech. This means that XTTS can adapt its speech output based on the situation, making it particularly useful for applications where the tone and style of speech are important. Researchers have also been combining large language models (LLMs) with TTS systems to create conversational agents that sound more human-like. For example, when ChatGPT is integrated with advanced TTS models, it can engage users in natural-sounding conversations, offering more personalized and interactive experiences. This technology has important applications in personalized virtual assistants and interactive media, where a more tailored voice experience can enhance user interaction.
@@ III-A Selection of Large Language Model (LLM)
Llama2-7BLlama3-8BGemma2-9BMixtral-7B001,00010001{,}0001 , $0002{,}$ $0003{,}$ 000716.84716.84716.84716.84530.4530.4530.4530.4579.13579.13579.13579.133,145.893145.893{,}145.893 , 145.89LLMTime Taken (seconds)
Figure 2: Time taken by different LLMs
We conducted an experiment to evaluate the performance of various large language models (LLMs) for our work. The experiment involved having the LLMs generate 50 conversation scenarios involving multiple participants, ranging from 2 to 5 individuals. We employed a checksum program to verify that the generated responses adhered to the required format. Following this, we recorded the time taken for response generation and the number of responses that deviated from the correct format. As described in Fig.2, our findings indicate that Llama3 generates responses more quickly than the other LLMs. Given that our work may require the generation of hundreds or thousands of responses, resource efficiency and the time taken per generation are critical factors for our work.
Llama2-7BLlama3-8BGemma2-9BMixtral-7B005555101010101515151520202020252525252020202022220012121212LLMNumber of wrong format
Figure 3: Number of responses in the incorrect format
Based on our second measurement in Fig.3, Gemma2 produced zero responses in the wrong format, while Llama3 had only two incorrect responses. After a comprehensive evaluation, we selected the Llama-3 8B model for conversation generation due to its fast speed and acceptable error rate, ensuring the generation of high-quality and diverse synthetic conversations.
@@ III-D Text-to-Speech Conversion
In this step, the generated text dialogues are converted into audio, with unique voices assigned to each speaker. For text-to-speech conversion, we utilize two different models for distinct purposes. The first model we employ is Parler-TTS, a text-to-speech synthesis model that generates natural-sounding speech from text, delivering high-quality and customizable audio outputs efficiently. Initially, we attempted to use Parler-TTS exclusively for generating the conversation audio. However, Parler-TTS has a limitation: it tends to produce variations in voices even when the same configuration is applied. As a result, the voice of a single persona, such as Alice, may change from one dialogue to another within the same conversation. To address this issue, we incorporate a second model, XTTS, which is a zero-shot text-to-speech model with voice cloning capabilities. Given an audio sample and text input, XTTS can clone the voice from the audio and generate speech using that cloned voice. However, XTTS is limited in its ability to produce multiple unique voices, aside from the few built-in demo voices. To achieve the desired audio quality and consistency, we combine the strengths of both Parler-TTS and XTTS. First, we use Parler-TTS to generate a unique voice for each persona. Next, we employ XTTS to clone these unique voices and generate the speech for each dialogue. The process for generating synthetic conversation audio is illustrated in Fig.7. This approach ensures that each persona maintains a consistent voice throughout the entire conversation and across the dataset.
@@ III-E Audio dialogues concatenation
Filename Start End Speaker
@@
TABLE I: Dialogue segments with start and end times.
The resulting audio files from the previous text-to-speech conversion step are labeled according to the order of their dialogues, such as Alice_0.wav, David_1.wav, and Alice_3.wav. These dialogues are then concatenated to form a complete conversation audio file, which is saved in .wav format. Following this, a ground truth CSV file is created, which includes the filename, starting timestamp, ending timestamp, and speaker information, as illustrated in Table I. Additionally, the conversation audio can undergo further post-processing, such as adding background noise or reverb.
@@ IV EXPERIMENT
In this section, we detail the experimental procedures and setups used to evaluate the effectiveness of the synthetic audio conversation generation framework. The experiment was designed to assess the quality and the validity of the generated conversations.
@@ IV-A Experimental Setup
The experiment involved generating a total of 200 synthetic conversation audios using the framework outlined in the Methodology section. These conversations varied in length and complexity, with the number of speakers per conversation ranging from 2 to 5. The personas used in these conversations were selected randomly from a predefined pool of 9 personas, ensuring a diverse representation of dialogue styles and topics. For the text generation phase, the Llama3-8B model as we mentioned earlier was utilized to generate coherent and contextually relevant conversations. The generated texts were then converted into audios using Parler-TTS for initial voice synthesis and XTTS for voice cloning to ensure consistency across dialogues.
@@ V-A Text Generation Time
During the text generation phase, the Llama3-8B model generated 200 conversation scenarios, involving 2 to 5 participants each. Out of the 200 generated conversations, 189 adhered to the correct format, resulting in a success rate of 94.5%. The total time taken for this process was 3730.06 seconds, averaging approximately 18.65 seconds per conversation. This efficiency demonstrates the Llama3-8B model’s suitability for large-scale synthetic data generation.
@@ V-B Text-to-Speech Conversion Time
The process of converting the generated text conversations into audio using Parler-TTS and XTTS took 4457.94 seconds in total, averaging approximately 22.29 seconds per audio file. The cumulative duration to generate 189 audio conversations was approximately 8,200 seconds in total, and the time required remains within acceptable limits for practical applications.
@@ V-C Exploratory Data Analysis
[…]
Figure 8: Distribution of Total Number of Speakers in Segments
The final dataset, consisting of 189 conversation audio files has a total duration of 4.01 hours. It is of high quality and suitable for various applications, including audio classification, speech recognition, and training voice assistants. The inclusion of ground truth annotations further enhances the usability of the dataset for supervised learning tasks. The analysis of the distribution of speaker appearances in the conversations reveals significant insights into the frequency and prominence of different speakers. As illustrated in Fig.8, the distribution of the total number of speakers in each segment varies, with segments involving 2 to 5 speakers. The pie chart shows that 2-speaker segments comprise 27.5% of the total, while 3-speaker, 4-speaker, and 5-speaker segments represent 22.2%, 25.9%, and 24.3%, respectively. This distribution indicates a relatively balanced participation across segments, with a slight tendency toward conversations involving fewer speakers.
0050505050100100100100150150150150200200200200250250250250300300300300350350350350AliceBenCathyEvaDavidHenryIsabellaGraceFrank300300300300278278278278246246246246238238238238234234234234229229229229215215215215206206206206205205205205Appearance count
Figure 9: Distribution of Speakers Appearance
Furthermore, Fig.9 presents a horizontal bar chart that details the individual appearances of each speaker. The analysis reveals that ”Alice” has the highest count at 300 appearances, followed by ”Ben” with 278 appearances, and ”Cathy” with 246 appearances. On the lower end, ”Frank” has the least number of appearances at 205, closely followed by ”Grace” with 206. The relatively even distribution among speakers suggests that the conversation segments are designed to give equal speaking opportunities to most participants, with some variability.
@@ VII FUTURE WORK
Based on the promising results of this experiment, several future research and development can be proposed.
VII-A Incorporating Environmental Contexts
Future research could explore the incorporation of synthetic environmental sounds or background noise to simulate more realistic conversation scenarios. This would provide additional context for the generated dialogues, making them more applicable to real-world audio classification and speech recognition tasks.
VII-B Evaluating Generalization to Other Languages
While this study focused on English, evaluating the framework’s ability to generalize to other languages is a critical area for future work. This would involve training the LLMs and TTS systems on multilingual datasets and assessing the quality and coherence of the generated conversations in different languages.
By addressing these areas, the framework could be further enhanced to meet the growing demands of synthetic audio generation and its applications in machine learning and artificial intelligence.

### v3

@@ I Introduction
The rise of large language models (LLMs) has transformed natural language processing, enabling coherent, contextually relevant text generation. This progress has improved advancements in various fields, including dialogue systems and automated content creation. However, one under-explored application is generating synthetic audio datasets for multi-speaker conversations, which are essential for training speech-related models. Existing audio conversation datasets are limited in size and diversity, hindering the development of robust models that handle diverse scenarios. To address this, we propose ConversaSynth, a framework for generating synthetic audio conversations using LLMs and text-to-speech (TTS) systems. ConversaSynth produces dialogues with distinct personas covering a wide range of topics, converting them into audio to create a diverse and realistic dataset. This resource supports applications such as:
  • Audio Classification: Enhancing the accuracy of models in identifying and categorizing different types of audio content.
  • Speech Recognition and Extraction: Improving models’ ability to transcribe and segment multi-speaker conversations.
ConversaSynth also allows for the creation of specific conversational topics and speaker attributes, filling gaps in existing datasets. This flexibility is vital for training models that perform well across varied contexts and speaker traits. Overall, ConversaSynth addresses current limitations and advances synthetic audio generation for future research and development in machine learning.
@@ II Literature Review
The integration of LLMs and TTS technologies has the potential to revolutionize applications such as virtual assistants, content creation, and accessibility tools. This review explores key developments in LLMs and TTS, along with their integration.
@@ II-A Large Language Models (LLMs)
LLMs have significantly advanced NLP by leveraging massive datasets to generate human-like text. Progress from GPT-2 to GPT-4 has led to notable improvements in understanding context and generating coherent text [1]. The Transformer architecture, introduced by Vaswani in 2017 [2], employs self-attention mechanisms to capture long-range dependencies, enabling more sophisticated text generation. OpenAI’s GPT-3, with 175 billion parameters, set new benchmarks for LLMs by excelling at various language tasks, and GPT-4 further refined these capabilities. LLMs now power applications like writing assistance, storytelling, and chatbots, demonstrating their versatility in generating high-quality content.
@@ II-B Text-to-Speech (TTS) Technologies
TTS technology has evolved from concatenative synthesis, which used pre-recorded speech segments, to deep learning models that generate more natural speech. Tacotron [3] and WaveNet [4] transformed TTS by using mel-spectrograms and neural vocoders, resulting in more realistic synthetic speech. Advanced models like Parler-TTS [5] and XTTS [6] further improved expressiveness and customization, adapting speech output to various contexts. Integrating LLMs with TTS systems, such as using ChatGPT with advanced TTS, has enabled more human-like conversational agents, enhancing personalized virtual assistants and interactive media.
@@ III-A Selection of Large Language Model (LLM)
TABLE I: Time Taken by Different LLMs
LLM Time Taken (seconds)
Llama2-7B 716.84
Llama3-8B 530.40
Gemma2-9B 579.13
Mixtral-7B 3,145.89
We conducted an experiment to evaluate the performance of various large language models (LLMs) for our work. The experiment involved having the LLMs generate 50 conversation scenarios involving multiple participants, ranging from 2 to 5 individuals. We employed a checksum program to verify that the generated responses adhered to the required format. Following this, we recorded the time taken for response generation and the number of responses that deviated from the correct format. As described in Table I, our findings indicate that Llama3 generates responses more quickly than the other LLMs. Given that our work may require the generation of hundreds or thousands of responses, resource efficiency and the time taken per generation are critical factors for our work.
TABLE II: Number of Responses in Incorrect Format by Different LLMs
LLM Number of Wrong Format Responses
Llama2-7B 20
Llama3-8B 2
Gemma2-9B 0
Mixtral-7B 12
Based on our second measurement in Table II, Gemma2 produced zero responses in the wrong format, while Llama3 had only two incorrect responses. After a comprehensive evaluation, we selected the Llama-3 8B model for conversation generation due to its fast speed and acceptable error rate, ensuring the generation of high-quality and diverse synthetic conversations.
@@ III-D Text-to-Speech Conversion
In this step, the generated text dialogues are converted into audio, with unique voices assigned to each speaker. For text-to-speech conversion, we utilize two different models for distinct purposes. The first model we employ is Parler-TTS, a text-to-speech synthesis model that generates natural-sounding speech from text, delivering high-quality and customizable audio outputs efficiently. Initially, we attempted to use Parler-TTS exclusively for generating the conversation audio. However, Parler-TTS has a limitation: it tends to produce variations in voices even when the same configuration is applied. As a result, the voice of a single persona, such as Alice, may change from one dialogue to another within the same conversation. To address this issue, we incorporate a second model, XTTS, which is a zero-shot text-to-speech model with voice cloning capabilities. Given an audio sample and text input, XTTS can clone the voice from the audio and generate speech using that cloned voice. However, XTTS is limited in its ability to produce multiple unique voices, aside from the few built-in demo voices. To achieve the desired audio quality and consistency, we combine the strengths of both Parler-TTS and XTTS. First, we use Parler-TTS to generate a unique voice for each persona. Next, we employ XTTS to clone these unique voices and generate the speech for each dialogue. The pipeline for generating synthetic conversation audio is illustrated in Fig.5. This approach ensures that each persona maintains a consistent voice throughout the entire conversation and across the dataset.
@@ III-E Audio Dialogues Concatenation
TABLE III: Dialogue Segments with Start and End Times
Filename Start (s) End (s) Speaker
@@
The resulting audio files from the previous text-to-speech conversion step are labeled according to the order of their dialogues, such as Alice_0.wav, David_1.wav, and Alice_3.wav. These dialogues are then concatenated to form a complete conversation audio file, which is saved in .wav format. Following this, a ground truth CSV file is created, which includes the filename, starting timestamp, ending timestamp, and speaker information, as illustrated in Table III. Additionally, the conversation audio can undergo further post-processing, such as adding background noise or reverb.
@@ IV EXPERIMENT
In this section, we show the experimental procedures and setups used to evaluate the effectiveness of the synthetic audio conversation generation framework. The experiment was designed to assess the quality and the validity of the generated conversations.
@@ IV-A Experimental Setup
We generated 200 synthetic audio conversations with 2 to 5 speakers using Llama3-8B for text generation and Parler-TTS combined with XTTS for voice synthesis and cloning. Nine personas (Alice, Ben, Cathy, Eva, David, Henry, Isabella, Grace, and Frank) were randomly assigned to ensure diversity in dialogue styles and topics. All experiments were conducted on an Nvidia GTX 1080Ti to ensure optimal performance.
@@ V-A Text Generation Time
During the text generation phase, the Llama3-8B model generated 200 conversation scenarios, involving 2 to 5 participants each. Out of the 200 generated conversations, 189 adhered to the correct format, resulting in a success rate of 94.5%. The total time taken for this process was 3,730.06 seconds, averaging approximately 18.65 seconds per conversation.
@@ V-B Text-to-Speech Conversion Time
The process of converting the generated text conversations into audio using Parler-TTS and XTTS took 4,457.94 seconds in total, averaging approximately 22.29 seconds per audio file. The cumulative duration to generate 189 audio conversations was approximately 8,200 seconds in total, and the time required remains within acceptable limits for practical applications.
@@ V-C Exploratory Data Analysis
The dataset includes 189 conversations totaling 4.01 hours, suitable for tasks like audio classification and speech recognition. Speaker distribution is fairly balanced, with 2-speaker segments making up 27.5%, and 3-, 4-, and 5-speaker segments comprising 22.2%, 25.9%, and 24.3%, respectively, as shown in Table IV.
TABLE IV: Distribution of Speakers per Conversation
Number of Speakers Percentage (%)
2 Speakers 27.5
3 Speakers 22.2
4 Speakers 25.9
5 Speakers 24.3
In terms of speaker appearances, Table V shows ”Alice” had the most with 300 appearances, while ”Frank” had the fewest at 205.
TABLE V: Speaker Appearance Distribution
Speaker Appearance Count
Alice 300
Ben 278
Cathy 246
Eva 238
David 234
Henry 229
Isabella 215
Grace 206
Frank 205
V-D Audio Quality Evaluation
To evaluate the quality of the generated synthetic audio conversations, we calculated the Signal-to-Noise Ratio (SNR) for each audio file. SNR is a commonly used metric in audio quality assessment, measuring the level of the desired signal relative to the noise. A higher SNR indicates better audio quality. In our evaluation using SNR, the average SNR across all evaluated audio files was found to be 93.49 dB, indicating a high level of audio clarity. This result suggests that the synthetic audio generated by our framework maintains excellent fidelity, enhancing its potential for practical applications in various domains.
@@ VII FUTURE WORK
Based on the positive results of this study, there are several possible directions for future research. Future work could integrate synthetic environmental sounds or background noise to create more realistic conversation scenarios, enhancing the applicability of the generated dialogues for audio classification and speech recognition tasks. Additionally, while this study focused on English, assessing the framework’s generalization to other languages is essential. This would involve training LLMs and TTS systems on multilingual datasets to evaluate the quality and coherence of conversations in diverse languages. By addressing these areas, the framework can better meet the growing demands of synthetic audio generation in machine learning and artificial intelligence.